OpenStudy (anonymous):

Distance = rate * time In the equation D = r * t, to solve for "r", one should: A.) Subtract "t" from both sides of the equation. B.) Divide both sides of the equation by "t". C.) Multiply both sides of the equation by "t". D.) Add a negative "t" to both sides of the equation.

OpenStudy (texaschic101):

D = RT D/T = RT/T D/T = R Divide both sides of the equation by T :)